Could you please help me to solve this exercise, i know it is quite long, but please do as much as as you are allowed to do. I appreciate very much if you can explain in details how the numbers come from so that I can apply using the way for similar exercises

Many thanks

) Cost assisgnment Evercise NYz manufactures two products: A and the following per-unit information A rview of the company accounting records revealed per-unit intformation Production volume (units) Direct material Direct labor 2,500 5,000 s 60 hours 2 hours Direct labor costs budgeted for the period are $ 240.000. The total amount of direct labor hours available is 24.000, Manufacturing overhead per units is currently computed by spreading overhead of $1,860,000 over direct labor hours. The variable portion of manufacturing overhead is $300,000. Management is considering a shit to activity-based accounting procedures, and the tollowing data are available costing in an effort to improve the firms Cost 240,000 1,500,000 120,000 1,860,000 Cost Driver Number of setups Cost Pool Setups General factory Machine processing 100 Direct labor hours5,000 2,200 20 15,000 800 Machine hours XYZ determines selling prices by adding 40% to a products total cost. Required: 1. Compute the per-unit cost and selling price of product B by using XYZs current costing 2. Compute the per-unit overhead cost of product B if the company switches to activity-based 3. procedures. costing Compute the total per-unit cost and selling price under activity-based costing.
Answer #1

1) Manufacturing overheas= total overhead costs/total direct labour hrs
=1860000/24000
=77.50
Total per unit cost of B=direct material+direct labor+mfg overhead
direct labour cost per hour=total costs/total hrs
=240000/24000=10
=60+(3*10)+(3*77.5)
=322.50
Marup=322.5*40%=129
selling price=322.5+129=451.50

2)Per unit overhead cost of B:
per setup cost=total setups cost/no of setup=240000/(100+20)=2000
genral factory cost per driver= total cost/direct labr hrs=1500000/(5000+15000)=75
Machine processing cost driver=total cost/total machine hrs=120000/(2200+800)=40
per unit overhead cost of B=(2000*20)+(75*15000)+(40*800))/total units=1197000/5000
=239.40


3)per unit cost of A= direct material+direct labor+mfg overhead
mfg overhead=(2000*100)+(75*5000)+(40*2200)=663000
per unit overhead cost=663000/5000=132.60
per unit cost of A=40+(2*10)+132.60=192.60
selling price=192.60*(1+markup)
=192.60*(1+40%)=269.64

per unit cost of B=60+(3*10)+239.40=329.40
selling price=192.60*(1+markup)
=329.40*(1+40%)=461.16